1. Find the force required to make a 50 kg mass of metal with a S.G. of 5.5, when it is half submerged in a. Water b. Oil with S.G. of 0.68.
2. Find the weight of a ball of volume 80 cm3 if it floats, half submerged on a liquid of density 1.2 g/cm3
1
Expert's answer
2020-03-16T13:05:19-0400

1)

a)

"W=F+0.5\rho gV"

"F=50(9.8)-0.5\frac{50(9.8)}{5.5}=445\ N"

b)


"F=50(9.8)-0.5\frac{50(0.68)(9.8)}{5.5}=460\ N"

2)


"W=0.5\rho gV"

"W=0.5(1200)(9.8)(80\cdot 10^{-6})=0.47\ N"